OpenStudy (anonymous):

Find the IVP by using Laplace Transforms y''+5y'-6y=21e^{t-1}; y(1)=-1, y'(1)=9

OpenStudy (amistre64):

well, take the laplace ... what do we get?

OpenStudy (anonymous):

Well i'm not sure if i need a translation on the t-axis? Taking the Laplace of both sides is fairly easy, but given the fact that the initial conditions are not at zero, i'm not sure what to do at the moment.

OpenStudy (amistre64):

yeah, thats a rare twist ... let r = t-1 if you want to substitute it

OpenStudy (amistre64):

im not that familiar with what a convolutionis so im not sure if that would be useful here

OpenStudy (amistre64):

y(t-1) = y(r) .... y(t=1) = y(r=0) is what im thinking

OpenStudy (anonymous):

If L{f(t)} exists for s>c, then L{u(t-a)f(t-a)}= e^{-as}F(s) for s>c+a

OpenStudy (anonymous):

I think this may be it

OpenStudy (amistre64):

by definition \[L\{y''(t-1)\}=\int_{0}^{\infty}y''(t-1)e^{-st}~dt \] let t = r+1, dt = dr \[L\{y''(r+1-1)\}=\int_{0}^{\infty}y''(r+1-1)e^{-s(r+1)}~dr \] \[L\{y''(r)\}=e\int_{0}^{\infty}y''(r)e^{-sr}~dr \]

OpenStudy (amistre64):

your find may be it, im just not that adept at all the transforms since i dont get to use it that often :)

OpenStudy (amistre64):

ive made an error in my thought ... e^(-sr) e^(-s) so e^(-s) pulls out as a constant

OpenStudy (anonymous):

your method looks adequate; you're using convolution. I may try that method.

OpenStudy (amistre64):

since i need the practice :) which means this could get messed up lol \[L(y'')+5L(y')-6L(y)=21L(e^{r})\] \[L(y'')+5L(y')-6L(y)=21L(e^{r})\] \[s^2L(y)-sy(0)-y'(0)+5(sL(y)-y(0))-6L(y)=21L(e^{r})\] \[s^2L(y)-sy(0)-y'(0)+5sL(y)-5y(0)-6L(y)=21L(e^{r})\] \[s^2L(y)+s-9+5sL(y)+5-6L(y)=21L(e^{r})\] \[L(y)(s^2+5s-6)=21L(e^{r})+4-s\] \[L(y)=\frac{25-s}{(s+6)(s-1)^2}\]

OpenStudy (amistre64):

21/(s-1) +4 - s so i messed up the last 2 rows i believe but its fixable lol

OpenStudy (amistre64):

17 +5s -s^2 might be the numerator ... if you want to let me know when and where i mess this totally up feel free :)

OpenStudy (anonymous):

Sounds good, ill let you know my answer in a bit.

OpenStudy (amistre64):

\[17 +5s -s^2=A(s-1)^2+(Bs+C)(s+6)\] \[17 +5(-6) -36=A(-6-1)^2\] \[-49=49A~:~A=1\] \[17 +5s -s^2=(s-1)^2+(Bs+C)(s+6)\] \[17 +5 -1=(B+C)(1+6)\] \[3=B+C\] \[17=(0-1)^2+C(0+6)\] \[8/3=C~:~B=1/3\] \[L(y)=\frac{1}{s+6}+\frac13\frac{s}{(s-1)^2}+\frac83\frac{1}{(s-1)^2}\] \[y(r)=e^{-6r}+\frac13 e^r(r+1)+\frac83e^rr\] \[y(t-1)=e^{-6(t-1)}+\frac13 e^{t-1}t+\frac83e^{t-1}(t-1)\] -------------------------------------- the wolf gives us a solution as: \[y(t) = e^{-6t-1} (3 e^{7t} (t-1)-e^7)\] \[y(t) = 3e^{6t-1} (t-1)-e^{-6t+6}\] so im pretty sure ive messed something up along the way, of course the wolf solved for y(t) which would be a shift of our y(t-1) maybe?

OpenStudy (anonymous):

wow! This one is brutal.

OpenStudy (amistre64):

yeah, and i think my decomp is messed up ...i hate typing and trying to math

OpenStudy (amistre64):

A = -1 not 1 messed up the rest of the process

OpenStudy (amistre64):

C=3

OpenStudy (amistre64):

B=0

OpenStudy (amistre64):

\[L(y)=3\frac{1}{(s-1)^2}-\frac{1}{s+6}\] \[y(r)=3r~e^r-e^{-6r}\] i wonder if i should put back on the e^-s, when i do the wolf gives me some heaviside function

OpenStudy (amistre64):

the wolf agrees with \[y(r) = 3r e^r-e^{-6r}\] :)

OpenStudy (anonymous):

I think they are using Duhamel's Principle. They're solving for the general solution then

OpenStudy (anonymous):

applying convolution

OpenStudy (anonymous):

good job.

OpenStudy (anonymous):

And thanks for the help.

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

Likewise amistre64. It was nice talking to you. Cheerio!